STAR Integrated Calibration/Validation System website

GOES ABI Calibration Events Log

This is a collection of events that, at the discretion of the GOES-R Calibration Working Group (CWG), have known or potential impact on L1b data quality. It is maintained by the CWG throughout the instrument mission life for ABI L1b users, producers, calibration specialists, managers, vendors, international partners, and members of the public interested in ABI calibration. For other ABI related events, please check the link below:

https://www.ospo.noaa.gov/Operations/messages.html

(Click column headings to sort; Type in the "Search" box to query table contents.)

Updated: 1/3/2022

Event StartDate EndDate Notes
B16 LUT update 05/10/2019 05/10/2019 The lunar rejection threshold of Band 16 was updated to reduce the potential striping caused by detector noise.
Navigation parameters update 6/1/2017 6/15/2017 RDP values are updated
Intermittent Channel 8 striping 12/31/2018 01/31/2019 repeated striping due to anomalous performance of det #208 and #107. Corrected with the BDS update of det #208 and manual space look reset of #107
ABI Band02 is ~7% brighter than VIIRS 6/1/2017 G16 ABI Band2 (0.64um) is brighter than VIIR I1band by ~6-8%, depending on time.
B05 Striping associated with Detector #200/#201 6/1/2017 6/24/2019 Persist striping patterns are detected associated with bad performance from Dectector #200/#201
B10 D312 Striping and BDS Update 6/1/2017 10/3/2017 Striping was caused by B10 Detector #312 and was removed with the BDS update on 10/03/2017
B12 D257 Striping and BDS Update 6/1/2017 10/3/2017 Striping was caused by B12 Detector #257 and was removed with the BDS update on 10/03/2017
Cold pixels around fire 6/1/2017 4/23/2019 For the 3.9um channel (B07, and for this band only), some pixels around (usually not immediately next to) extremely hot targets such as fire have substatially cold bias. The affected pixels are properly flagged. Corrected onon 04/23/2019. Major navigation upgrade,6/1/2017,,Upgrade elements: (i) static FPA shifts are absorbed into the LUTs; (ii) GNC gyro rate capping is adjusted. Upgrade effects: (i) navigation disturbances from the Kalman filter reset are reduced several times; (ii) navigation stability is increased. Periodic IR Calibration Anomaly (PICA),6/1/2017,10/19/2017,PICA was a periodic radiometric calibration anomaly with 15 minutes of periodicity observed at GOES-16 ABI Mode-3 MESO and CONUS images caused by the residual of incident angle-dependent scan mirror emissivity corrections. It was corrected with the update of scan-mirror emissivity LUTs on 10/19/2017. Sub-optimal solar calibration for VNIR Channels,01/15/2017,12/12/2017,Until 12 December 2017
ICT Temperature Bias 6/1/2017 6/19/2018 The ICT PRT coefficients that converted the PRT raw counts into temperatures were inconsistent with the values from the correct vendor source. This discrepancy led to a -0.2 K bias to the calculated ICT temperatures and colder bias of similar magnitude of the L1b products for all the IR channels. On 19 June 2018, the revised LUT was implemented in GS data processing and resulted in improved radiometric calibration accuracy.
GOES-16 Drifted from 89.5W to 89.2W 8/8/2017 9/3/2017 GOES-16 drifted from 89.5W to 89.2W for the purpose of testing potential impacts related to shifting GOES-16 operational station keeping position from 75.0W to 85.2W.
Transient navigation anomalies 9/28/2017 Anomalies involve large and sudden East-West and/or North-South navigation swings exceeding MRD limits. The reason identified as an intermittent loss of the gyro rate information in the Ground processing. The fix extrapolates observed gyro rate onto the period of absence.
B05 Striping caused by incorrect of use of Look-up table 11/19/2017 12/29/2017 In-correct solar calibration LUT (pre-launch solar calibration coefficient) was used, resulting stronger striping at B05 during this period
Major navigation upgrade 11/20/2017 11/22/2017 Upgrade elements: (i) RDP correction for rotation and stretch; (ii) diagonal scan is implemented for 064 star detection; (iii) dynamic LMC is implemented; (iv) Kalman filter retuned. Upgrade effects: significant reduction of all MRD-specified navigation errors to well-below the specs. Deficiency: Kalman filter-induced instability remains in the navigation of thermal channels, though the magnitude is significantly reduced.
GOES-16 Drifted from 89.3W to 75.2W 11/29/2017 12/18/2017 GOES-16 station keeping relocated from the 89.3W Test location to its 75.2W operational location,and intruments re-evaluted for operational readiness.
B05 10% jump 12/12/2017 12/29/2017 In-correct B05 solar calibration coefficient was used during this period GOES-16 ABI L1B data operational
B12 D290/291 Striping and BDS update 12/30/2017 1/2/2018 Stripings have been detected for B12 D290 & D291 and the corresponding stripes were removed with the BDS update on 01/02/2018
VNIR Operational Algorithm Update 1/17/2018 2/7/2018 A new calibration algorithm was operationally implemented to reduce the VNIR striping observed since in-orbit. This implementation was accomplished with two steps: the first step was the upload of the new detector non-linearity (Q) LUT for B01/B02/B03 on 01/17/2018 and the second step was to implement the new algorith using Q-scaling on 02/07/2018. This new algorithm resulted in significant reduce in striping at B01/B02/B03 and slight change in the VNIR radiometric calibration accuracy. No change in B05 striping.
Sub-optimal solar calibration for VNIR channels 2/14/2018 2/20/2018 The three extra solar calirbation events were conducted to simulate the similar illumination condition for the solar calibration event conducted at about one year ago
B12 D309 Striping and BDS update 2/26/2018 3/5/2018 Striping was caused by B12 Detector #309 and was removed with the BDS update on 03/05/2018
Solar Intrusion 3/2/2018 Strong stray light appearing mostly in the Southern hemisphere. Two possible reasons: (i) Earth exclusion zone was not updated for the actual GOES-EAST position at -75.2 degrees; (ii) atmospheric refraction was not taken into account in the calculations of the Earth exclusion zone
Super Caterpillar Tracks 3/26/2018 3/26/2018 SCT correlates with the scan angle deviations, encoder spike flags, and enhanced high-energy electron flux in SEISS.
Sub-optimal solar calibration for VNIR channels 3/27/2018 3/27/2018 This extra solar calibration was conducted to simulate the similar illumination condition for the solar calibration event at about one year ago
VNIR 10% jumps 4/10/2018 4/10/2018 All 6 VNIR channels experienced ~10% jumps in the radioemtric calirbation accuracy due to the incorrect use of solar calibration coefficients
Sub-optimal solar calibration for VNIR channels 4/24/2018 4/24/2018 This extra solar calibration was conducted to simulate the similar illumination condition for the solar calibration event at about one year ago
Earth Orientation Parameter Predictions (EOPP) Correction 4/27/2018 Time format mismatch between GPS and ABI is identified and removed. In the past, that led to a small navigation bias for all channels. It was temporarily compensated since June 2018 by a constant correction. However, the bias is time dependent, so only now it is permanently fixed.
Channel 8 striping 4/29/2018 4/29/2018 Channel 8 striping was discovered. Caused by detector response level shifts (observed by both SPL and ICT data) with Detector #286 (Indexing from zero). It was discovered that the space look (SPL) used to calculate the operational calibration coefficients was locked during this period
Channel 10 striping and BDS update 5/19/2018 5/22/2018 Channel 10 striping was discovered. Caused by ground processing spacelook latchup after Detector #294 (indexing from zero) response level shifts (observed by both SPL and ICT data). Corrected with a BDS change.
Channel 10 Striping 6/3/2018 6/4/2018 Channel 10 striping was discovered. Caused by ground processing spacelook latchup after Detector #156 (indexing from zero) response level shifts (observed by both SPL and ICT data). This led to a bias of affected pixels that appears as a strip in image. Corrected by resetting space look count.
Channel 9 Striping 6/4/2018 6/4/2018 Channel 9 striping was discovered. Caused by ground processing spacelook latchup after Detector #303 (indexing from zero) response level shifts (observed by both SPL and ICT data). This led to a bias of affected pixels that appears as a strip in image. Corrected by resetting space look count.
Transient navigation anomaly fixed 10/15/2018 10/15/2018 Large navigation errors in all channels lasting for about one hour with a random occurrence were associated with the gaps in the Ground gyro rate processing and fixed in the Ground software upgrade.
Channel 9 striping and BDS update 10/27/2018 11/2/2018 Striping was discovered. Corrected with the BDS update.
Erroneous radiometric cal for VNIR bands 01/18/2019 01/22/2019 Incorrect VNIR bands gain began at 15:00UTC on 01/18/2019 and ended at 16:00UTC on 01/22/2019.Obvious striping was observed in bands 1 and 2 during the period.
Missing INST-CAL-M2 and EPH on 12/26/2018 12/26/2018 12/26/2018 Missing INST-CAL-M2 and EPH on 12/26/2018 from GS. However
DO.07 has a code bug in implemetation of CDRL 80 Rev F 10/15/2018 In equation 6.5-1 the integration factor was erroneously divided twice during the Q-scaling in DO.07.( WR5407/WR6348).
Mode 6 72-hour test run 2/19/2019 2/22/2019 Default Mode 6 version was run from 15 to 15 UTC.
GOES-16 HBT thruster Flush 3/12/2019 3/12/2019 A GOES-16 HBT Flush was performed from 2305UTC to 2315UTC with 10 minutes of product outage.Product navigation may be degraded for up to 45 minutes after the conclusion of the HBT flush. Event ended with a navigation anomaly - soft reset of Kalman filter elevating the image navigation errors for about three hours.
A persistent shift of BTD in GOES-16 6/15/2017 Bob Kuligowski group found a persistent shift in the time series of max and min bands 11-13 BTD values since 15 June 2017.
G16 band16 striping 03/31/2019 03/31/2019 Channel 16 detector #211 (0-based index) spacelook count mean was locked (latch-up) from 00:00z to 18:40z on 2019-03-31. Leading to more serious striping during the period. The latch-up was caused and released automatically with the random fluctuation of the spacelook.
G16 Transition to Mode 6 04/02/2019 New baseline Mode 6A is started at 16:00 UTC ; it replaced Mode 3E; link: https://www.ospo.noaa.gov/Operations/GOES/16/GOES-16%20Scan%20Mode%206.html
Unexpected G16 VNIR gain jumps 04/08/2019 04/09/2019 All the G16 VNIR bands experienced a unexpected large jump (>10%) after the GS software update (DO.07.02) on 04/08/2019. An unknown origin of the gains were used after this GS update. On 04/09/2019 the previous solar cal gain values derived with DO.07 using the 03/26 solar cal. data were re-applied to generate L1B data. Visible striping can be observed in the G16 L1B images during this abnormal period.
CPAF Update 04/23/2019 04/23/2019 Band 7 resampler updated at 20:15 UTC to correct the cold pixels around fire (CPAF).
G16 B02 bias correction 04/23/2019 1)GOES-16 new B02 K-LUT was implemented on GS at 11:00UTC on 04/23/2019(B02 with CWG's delivery. The other bands corrected with CDRL079_FM1_RevF); 2) calibration algorithm with RevG 3) instrument-to-spacecraft alignment angles with CDRL79-FM1_RevM. The major impact is that G16 B02 radiance is reduced by about ~6.2%.
North-South navigation anomaly 4/13/2019 04/13/2019 Anomalous North-South navigation shifts were observed by CWG in the full disk image at 23:00 UTC. The shift is strongest in the VNIR channels (up to 6 urad). The CONUS and MESO images produced in the period 23:00 to 23:10 may also be affected.
Navigation anomaly 05/04/2019 05/04/2019 Navigation anomaly and gap in data occurred from 12:00 to 13:00 UTC. IR channels may feature slightly elevated navigation errors (by less than 10 urad) during the next 24 hours.
Update the Kalman filter control LUTs 05/10/2019 05/10/2019 Deployed at 15:29 UTC this update is a temporary fix for the anomaly reported by ADR 955. This temporary fix will resolve the periodic Kalman filter issues that have been seen in the recent weeks and reported in ADR 955. The fix will be removed when the full fix is delivered in DO 09.
Update of Solar Avoidance Earth Mask 8/30/2018 8/30/2018 A new solar avoidance earth mask is implemented to: (1) Compensate for the difference as GOES-16 flying as GOES-East at 75.2W longitude while G16 ABI on-board configuration at 75.0W; and (2) Avoid of atmospheric refraction effects. Data outage occurred at 17:10Z - 17:15Z.
Possible image navigation problems 07/10/2019 07/10/2019 ABI images taken around 15:30 UTC may have navigation problems due to not fully-compensated moment adjustment.
Major upgrade of Product Generation (PG) at Ground System(GS) to DO.08.00.02 7/25/2019 including the correction of several issues concerning L1b products and the implementation of Predictive Calibration to improve GOES-17 Channel 8-16 calibration when the FPM temperature becomes unstable.
G16 ABI synchronization anomaly with PDA processing 08/17/2019 08/17/2019 For the period from 2019-08-17 0600 UTC to 2019-08-17 1650 UTC
G16 ABI synchronization anomaly with PDA processing 08/17/2019 08/18/2019 For the period from 2019-08-17 2120 UTC to 2019-08-18 0340 UTC
G16 band 9 stripes 08/26/2019 08/27/2019 G16 Channel 9 detector #31 (0-based index) spacelook count mean was locked (latch-up) at 22:30z on 2019-08-26 and caused stripes in L1b image after the time.The latch-up was released by manual spacelook reset at 12:26z on 2019-08-27 and the stripes disappear.
G16 ABI navigation anomaly 08/23/2019 08/23/2019 Navigation shift is detected by CENRAIS in the full-disk image of 20:30 UTC in the EW direction for all channels and for VNIR channels in NS direction. In the channel 2 image the shift is about 23 microradians in the East direction and 6 microradians in the South direction. Similar shifts may be present in the other imagery taken between 20:30 and 20:40 UTC.
G16 maneuver 10/30/2019 10/30/2019 ABI on GOES-16 had a scheduled data outage from 22:50 to 23:00 UTC due to the satellite maneuver. The images between 23:00 and 23:10 may have blank spaces
G16 navigation anomaly 11/01/2019 11/01/2019 ABI on GOES-16 had a data outage from 13:10 to 13:40 UTC due to antenna outage (intermediate frequency distribution switching switch error). During this period the images may have blank spaces and feature large navigation errors or be absent. First good full disk image is timed by 13:40 UTC.
G16 new Mode 6 took place in operation 11/22/2019 11/22/2019 The new G16 Mode 6 timeline took place in operation from 16:10 UTC.
GOES-16 L1b Data Gap 11/12/2019 11/12/2019 Based on ESPC GOES-R email notification G16 Product outage (ground failover) from 18:25-18:50 UTC for product delivered to PDA and GRB. G16 FD image at 1820. 1830 and 1840 UTC were missing.
GOES-16 L1b Data Gap 11/13/2019 11/13/2019 G16 FD image at 18:10 UTC was found missing and 18:20 image is incomplete due to the ground processing switching. That may also apply to the images in other formats.
GOES-16 L1b Data Gap 11/22/2019 11/22/2019 Due to the ABI Fine Track Error Calibration the GOES 16 ABI L1b data was missing for 10 minutes (1600-1610 UTC).
ABI FTE verification 11/25/2019 11/25/2019 Due to the ABI Fine Track Error Verification the GOES 16 ABI L1b full-disk images at 1700 and 1710 UTC were missing. That may also apply to the images in other formats.
L1b Data Gap 11/13/2019 11/13/2019 G16 FD image at 18:10 UTC was found missing and 18:20 image is incomplete due to the ground processing switching. That may also apply to the images in other formats.
Channel 10 Striping 12/04/2019 12/06/2019 G16 Channl 10 detector#221 (0-based index) spacelook count mean was locked (latch-up) at 16:00Z on 2019-12-04 and caused stripes in L1b images. The latch-up was released by mannual spacelook reset on 2019-12-05 and the stripes disappear. The BDS of this detector was updated at 15:05Z on 2019-12-06 that completed the resolution of this anomaly.
Channel 12 Striping 12/05/2019 12/18/2019 G16 Channel 12 detector #96 (0-based index) spacelook count mean was locked (latch-up) at 16:10Z on 2019-12-05 and caused stripes in L1b images. The latch-up was released by manual spacelook reset on the same day at 19:50Z and the stripes disappear temporarily. Repeated "latch-up" occurred several times later and caused stripes in L1b images. The BDS was updated on 12/18/2019 to resolve this issue.
Band 10 Striping 12/30/2019 12/31/2019 G16 band 10 detector#300 (0-based index) spacelook count mean was locked (latch-up) at 04:00Z and caused stripes in L1b images. The latch-up was released by the BDS update of the detector on 12/31 at ~16:00Z.
B08 striping and BDS update 03/03/2020 03/05/2020 G16 band 8 detector #148 (0-based index) BDS was updated at ~14:25UTC on 03/05/2020 to resolve the image striping caused by the unstable detector responsibility starting at ~10:28UTC on 03/03/2020.
Real-time op. instr. cal files outage 02/23/2020 The real-time operational instrument calibration files are not available via PDA.
G16 Navigation Anomaly 2/8/2020 2/8/2020 Elevated navigation errors was observed in all channels for the 20:50 UTC image. Anomaly is similar to the other events related to a gyro rate loss.
G16 Navigation Anomaly 2/10/2020 2/10/2020 Elevated navigation errors was observed in all IR channels for the 5:10 UTC image. Anomaly is similar to the other events related to a gyro rate loss.
GOES-16 INR anomaly 05/04/2020 05/04/2020 Large North-South image navigation errors in the 12:00 UTC FD images (other formats may also be affected close to that time) due to reset of the Kalman filter. Channel errors are up to 40 urad for IR and up to 12 urad for VNIR channels. Details are in the presentation by S. Guo and V. Kondratovich at the 2020-05-27 CWG meeting.
GOES-16 INR anomaly 05/28/2020 05/28/2020 Elevated East-West and North-South FD image navigation errors at 7:10 UTC due to a gyro rate gap. Channel errors are up to 15 urad. Other image formats between 7:00 and 7:10 UTC may also be affected. Details are in the presentation by V. Kondratovich and S. Guo at 2020-06-03 CWG meeting.
GOES-16 HBT Flush 06/23/2020 06/23/2020 HBT thruster flush was performed as scheduled at 23:00 UTC. No images from 23:00 to 23:10 except partial full-disk image at 23:00. Details are in the presentation by S. Guo and V. Kondratovich at the 2020-06-30 CWG meeting.
G16 ABI Reset Anomaly 07/26/2020 07/27/2020 Due to an instrument reset anomaly G16 ABI experienced an about six and half hours of L1B data outage from ~20:36Z on 07/26/2020 to 03:00z on 07/27/2020. Abundant stripings from weak to moderate magnitudes can be seen at Band16 images between 03:00- ~03:26z. It was followed by the striping caused by one single detector till ~04:24z. Details in the talk in the CWG meeting on 07/29/2020.
B10 striping and BDS update 07/28/2020 08/07/2020 The BDS of B10 Row#036 was updated at ~15:42z on 08/07/2020 to resolve the intermittent striping caused by the unstable detector starting at ~01:50z on 07/28/2020. More information is available at the status report in the 08/12/2020 CWG meeting.
B12 striping 09/17/2020 09/17/2020 Striping can be observed at G16 B12 images from 02:06UTC to 14:20UTC on 09/17/2020. This striping was caused by the spacelook latch-up. Details will be available in the CWG report on 09/23/2020.
B10 Striping & BDS update 09/17/2020 09/22/2020 Band10 images experienced striping caused by a noisy and unstable detector (Det#073). The issue was resolved with the BDS on 09/22/2020. Detailed information was reported in the CWG meeting on 09/23/2020.
B08 striping 09/05/2020 09/23/2020 Striping may be observed at B08 images from ~00:24 - 23:10z on 09/05/2020 and from 04:47 - 06:17z on 09/23/2020. The striping was caused by the latched spacelook of Det#273 during this period. Details can be found in the CWG meeting report on 09/09/2020 and 09/23/2020.
B10 striping 09/23/2020 09/23/2020 B10 images experienced a short-term striping between 14:40 - 14:50z on 09/23/2020. The striping was caused by the spacelook latch-up of Det#245 during this period.
G16 B01 solar cal data anomaly 09/17/2020 09/17/2020 After the solar calibration event on 09/17/2020 the G16 B01 images were black for 3.5 hours (11:00UTC to 14:30UTC) . This anomaly was caused by the erroneous B01 solar calibration data and was corrected by restoring the previous solar cal data conducted on 06/18/2020. Detailed information can be found in CWG meeting on 09/23/2020.
Band 01 gain update 11/04/2020 11/04/2020 A new set of GOES-16 Band 01 gains which were generated based on the solar calibration event on 09/17/2020 were applied at 22:29Z on Nov 4 2020. This is to correct the original solar calibration processing anomaly on 17 Sep 2020. Users may see a small shift (approximately 0.6%) in the B01 radiances due to this change.
B16 striping 11/21/2020 12/10/2020 The G16 B16 images experienced intermittent striping from 11/21/2020 at ~21:15 UTC to 12/10/2020 at ~22:19 UTC. The striping was caused by the unstable response of Det#211 during this period. References: CWG meetings on 12/02/2020 - 12/16/2020.
B09 striping 12/10/2020 12/10/2020 The G16 B09 images experienced striping from 12:10 UTC to ~17:30 UTC on 12/10/2020. This striping was caused by the spacelook latchup of Det#178. Reference: CWG meeting on 12/16/2020.
B02 stripes 12/15/2020 12/15/2020 Broad black stripes can be observed at the G16 B02 images from ~11:00Z to 15:47Z after the solar calibration event on 12/15/2020. Details at the CWG slides on 12/16/2020.
B10 striping 12/24/2020 12/24/2020 Striping at G16 B10 may be observed on 12/24/2020. It was caused by the space-look latch-up due to a sudden response change at det#094 at 10:43 UTC. The striping disappeared as the latched spacelook was naturally released at ~14:01 UTC. Details in the CWG meeting report on 01/06/2021.
B11 striping 12/31/2020 01/01/2021 Striping at G16 B11 may be observed from ~23:04 UTC on 12/31/2020 to ~15:40 UTC on 01/01/2021. This was caused by the latched space-look due to a sudden change of a detector response (det#234) at ~23:04 UTC on 12/31/2020. The striping disappeared as the latched spacelook was manually released at ~15:40 UTC on 01/01/2021. Details in the CWG meeting report on 01/06/2021.
B14 Striping and BDS update 12/22/2020 01/05/2021 Striping may be observed at B14 L1b images from 12/22/2020 and 01/05/2020. It was resolved with the BDS update for Det#240 on 01/05/2020. Details available in the CWG meeting on 01/06/2021
B08 Striping 01/05/2021 01/05/2021 B08 L1b images display a weak to moderate striping from ~06:58Z to ~14:14Z on 01/05/2021 due to the latch-up of the spacelook count for Det273. Details available in the CWG meeting on 01/06/2021.
ABI data outage 01/07/2021 01/07/2021 No ABI images were produced from 16:00 to 16:19 UTC due to the scheduled encoder calibration. Some blank FD images and partial MESO1 images were found for 16:00 UTC. No degradation was observed in the images produced outside of the period of calibration. Refer to CWG presentation by Song Guo and Vladimir Kondratovich on 1/13/2021.
ABI data outage 01/07/2021 01/07/2021 No ABI images were produced from 17:30 to 17:49 UTC due to the scheduled fine track error calibration. No degradation was observed in the images produced outside of the period of calibration. Refer to CWG presentation by Song Guo and Vladimir Kondratovich on 1/13/2021.
Degraded and missing images 01/19/2021 01/19/2021 Partial FD images at 17:30 and 17:40 UTC. Partial CONUS images for 17:31 UTC and no CONUS images for 17:36 UTC. Partial MESO1 images for 17:33-17:34 UTC. No MESO2 images for 17:33 UTC and partial MESO2 images for 17:34 UTC. Elevated navigation residuals at 17:50 UTC FD and CONUS images. Refer to Song Guo and Vladimir Kondratovich slides at the CWG meeting on 2021-01-27.
Image gap 01/21/2021 01/21/2021 Because of the work on flight software (FSW) update
Band15 striping 01/21/2021 01/22/2021 Weak to moderate striping can be observed in the G16 B15 images after the FSW update from 22:10 UTC on 01/21/2021 to 13:30 UTC on 01/22/2021. The striping was caused by the response level-shift induced spacelook latch-up. Details in the slides at the CWG meeting on 2021-01-27.
B07 striping 01/27/2021 01/27/2021 G16 B07 images experienced degraded images from 06:12Z to ~14:12Z on 01/27/2021. Root cause is likely due to the detectors' response to the space weather event. Details at the reports in the CWG meeting on 01/27/2021.
Bad Images Related to the Solar Cal Event 03/15/2021 03/15/2021 The GOES-16 all-band full-disk images at 11:10 UTC were incomplete as the result of the solar calibration event starting at 11:11 UTC. The B01 image was filled with fillvalue and B02 displayed broad black stripes at 11:15UTC. All the images returned to normal at 11:20 UTC. As part of the contingent plan to resolve the solar calibration anomalies occurred in 2020; there was no update of the solar calibration coefficients after this solar calibration event. Detailed information can be found in the CWG meeting on 03/17/2021.
VNIR channel gains update via BYOG 04/05/2021 04/05/2021 GOES-16 VNIR channel gains were implemented into operation via Bright-Your-Own-Gain (BYOG) at 17:51 UTC on 5 April 2021. The gain CWG delivered was based on the scheduled solar calibration event conducted on 03/15/2021. The effort was to minimize the impact of abnormal PG solar calibration processing. Detailed information can be found in the CWG meeting on 03/31/2021 and 04/07/2021.
B12 Stripings & one BDS update 04/19/2021 05/05/2021 Stripings at different locations can be observed at B12 images during this period. The BDS update for Det#188 was conducted on 05/05/2021. Details are available in the CWG meeting reports on 04/21/2021 ; 04/28/2021 and 05/05/2021.
B08 Striping 05/13/2021 05/13/2021 Striping appeared in the G16 B08 images between 08:16Z and 14:22Z on 05/13/2021. Details in the CWG presentation on 05/19/2021.
G16 data outage 05/05/2021 05/05/2021 BI on GOES-16 had a data outage from 20:50 to 21:39 UTC due to a spacecraft maintenance work. During this period all images are absent. The images around this gap have nominal navigation. Detailed discussion was presented at 2021-05-12 CWG meeting by S. Guo and V. Kondratovich.
B10 striping and BDS update 06/09/2021 06/10/2021 G16 B10 displayed some striping in the images collected on 06/09/2021 and 06/10/2021. A BDS event was conducted on 06/15/2021. More details available in the reports presented in the CWG meeting on 06/16/2021.
ABI image sequential anomaly 05/29/2021 05/29/2021 ABI L1b all channel images were not distributed between 20:59-21:35 UTC. Partial FD images were distributed at 21:30. Refer to ESPC Notification and S. Guo and V. Kondratovich CWG meeting presentation on 6/02/2021.
B09 striping and BDS update 06/14/2021 06/24/2021 G16 B09 images displayed intermittent striping since 06/14/2021. The issue was resolved with a BDS update on 06/24/2021. Details are available in the CWG meeting report on 06/23/2021.
B09 striping 07/16/2021 07/16/2021 B09 images displayed strong striping from 09:20 - 09:30Z on 07/16/2021. Root cause investigation was reported in the CWG meeting on 07/28/2021.
Striping in MESO2 images by a lunar intrusion 08/23/2021 08/23/2021 Apparent striping lines were observed in B08/09 MESO2 L1b images only at ~18:29 UTC on 08/23/2021. It was caused by a defect in the ELR algorithm. Detailed information is available in the CWG presentation on 09/22/2021.
Single thin stripe on B09 11/16/2021 11/16/2021 A thin stripe was present in a single swath due to lunar intrusion into the space look. The single pixel stripe is caused by a known issue in the enhanced lunar rejection algorithm. The contaminated spacelook observations occurred on B09 on 11/16 during the 1410Z timeline.
Single thin stripe by lunar intrusion 10/17/2021 10/18/2021 A thin single-pixel stripe was present in a single swath in two separate timelines caused by a known issue in ELR. The contaminated spacelook observations occurred at the following times for the specified bands and caused degradation in the products scanned during that time: 10/17 1410Z band 11 and 10/18 1440Z bands 10. Reference: CWG meeting slides by D. Goldberg on 10/27/2021.
Single thin stripe on B08 12/11/2021 12/11/2021 Description: A thin stripe was present in a single swath due to lunar intrusion into the space look. The single pixel stripe is caused by a known issue in the enhanced lunar rejection algorithm. The contaminated spacelook observations occurred on B08 on 12/11 during the 1040Z timeline (at 10:48:35Z). Details in the CWG meeting on 12/22/2021.
 
Event StartDate EndDate Notes
Detectors added to dead row list 12/20/2018 02/19/2019 The following detectors are added to the dead row list and skipped from l1b image generation: B8D91 B10D50 B10D149 B10D154 B11D35 B13D196 B15D333. This is a temporary method applied to reduce the strong striping at these IR bands.
Channels 8/10/11/12 striping 01/12/2019 01/18/2019 Striping due to elevated FPM temperature
G17 ABI Reset 1/3/2019 01/03/2019 G17 ABI was brought offline to implement a flight software update at ~1800UTC for two hours on 01/03/2019. GOES-17 ABI may experience degraded performance upon return to service while image navigation performance converges.
G17 ABI Mode6 test 12/27/2018 12/30/2018 Mode 6JKL timeline test
FPM peak temperature anomaly 11/25/2018 11/27/2018 Cryo-cooler reset tests during the "hot" FPM temperature periods on 11/25/2018-11/27/2018. Resulted in degraded IR imagery during these periods.
Mode 3G test run 11/15/2018 11/16/2018 Mode 3G test run (static spacelook); no FD L1B data during this period.
Detector map update 11/15/2018 11/15/2018 To improve the image navigation. Deployed at 18:48 UTC.
Kalman filter reset 11/15/2018 11/15/2018 Image navigation has temporarily degraded after KF reset at 18:32 UTC
B11 BDS update 11/14/2018 BDS for detector #203 updated
G17 drift to GOES-West 10/24/2018 11/13/2018 GOES-17 drifted from the test position (89.5W) to the GOES-West position (137.2W)
Data outage 10/23/2018 10/23/2018 L1B data outage between ~10:27 - ~13:02 UTC due to the flooding of sample outlier files (SOF) when scene saturated. The SOF clogged up the resampler and services crashed.
Detector map update 10/19/2018 10/19/2018 To improve the image navigation.
Visible star observation update 10/15/2018 10/15/2018 To improve precision of the star observation. Made in the framework of the ground processing upgrade.
Kalman filter parameter update 10/9/2018 10/9/2018 To improve convergence and stability of the filter. Translates to better image navigation.
Mode 3 timeline started 10/5/2018 Mode 3 started - Mode 6 ended.
Detector map update 10/4/2018 10/4/2018 To improve the image navigation. After update the Meso 1 and Meso 2 scans have used different detector maps for some time. The issue is under investigation.
Mode 6 started 10/2/2018 10/5/2018 Mode 6 started - Mode 4 ended.
Mode 4 started 10/1/2018 10/2/2018 Mode 4 started - Mode 6 ended.
Mode 6 started 9/26/2018 10/1/2018 Data distribution resumed with ABI in the Mode 6.
Yaw flip 9/24/2018 GOES-17 was conducted yaw-flip maneuver to inverted position.
Detector map update 9/20/2018 9/21/2018 To improve the image navigation.
Detector map update 9/13/2018 9/13/2018 To remove 3-pixel bias to the South in the LWIR channels. Deployed at 12:55 UTC.
BDS upload for IR bands 9/12/2018 BDS updates for large NEdT IR detectors.
Star detection parameter update 9/7/2018 9/7/2018 To improve the image navigation and stability. Deployed at 17:07 UTC.
Mode 3C implementation 9/5/2018 Mode 3C was deployed at 19:00 UTC. Results of the second and third blackbody calibrations only used for CONUS and MESO images.
Line Of Sight stability improvement 8/30/2018 8/30/2018 To improve the image navigation stability.
Band02 is ~9% brighter than SNPP/VIIRS 8/28/2018 G17 ABI B02 radiance is ~9% brighter than collocated SNPP VIIRS measurements.
PICA 8/28/2018 some of G17 ABI IR bands experience swath-to-swath meso-to-meso and CONUS-to-CONUS calibration variations
Star observation angle correction 8/9/2018 8/9/2018 To increase the number of observed stars.
Star observation angle correction 7/9/2018 7/9/2018 To enable IR star observation. That allowed to compensate for VNIR - MWIR offset. LWIR offset remained not fully compensated.
Mode 3 with extra stars started 7/5/2018 7/6/2018 Extra star observations increase stability of the image navigation.
Mode 3 started 7/2/2018 9/5/2018 Line Of Sight correction has enabled the start of the Mode 3 timeline.
Line of Sight correction 7/2/2018 7/2/2018 The correction has removed static offsets in the ABI pointing.
Yaw flip maneuver to original position 6/20/2018 6/20/2018 Performed between 17 and 19 UTC.
First Light Image 5/20/2018 5/20/2018 First Light image released.
Yaw flip maneuver to inverted position 5/16/2018 5/16/2018 Performed between 17 and 19 UTC.
Occurrence of "artificial" striping for the VNIR bands after DO7 10/15/2018 10/24/2018 Performed during this period; disappeared after drift.
Inaccurate GS solar gains on 10/18 and 11/24 10/18/2018 11/24/2018 Sudden gain drop in Band04.
Sudden change of solar cal gains at not-solar-calibration event time 11/14/2018 11/24/2018 Impact unknown on L1B images
Large oscillation of B02/B03 solar cal gains 8/30/2018 10/18/2018 Possible due to SD BRDF uncertainty.
VNIR bands incorrect gain 01/18/2019 01/19/2019 Incorrect VNIR bands gain began at 15:00UTC on 01/18/2019 and ended at 3:00UTC on 01/19/2019 after the solar calibration on 01/19/2019.
ABI hard reset 11/20/2018 11/21/2018 Began at 20:30 UTC and ended at 2:00 next day. Caused by the cryocooler anomaly. Large calibration and navigation anomalies afterward were caused by the need to reload various LUTs.
Crycoler anomaly 11/20/2018 Anomaly with the crycooler started around 10:15 UTC until the reset at 20:30 UTC
Crycooler anomaly 11/25/2018 Anomaly with the crycoller around 14:00 UTC; associated ABI reset also caused a range of calibration and navigation anomalies.
Mode 6F test run 11/27/2018 11/30/2018 Navigation is in family with the previous versions of Mode 6.
B09 BDS update 12/10/2018 Det#2:Column6->4. Det#52:Column6->1. Det#127:Column6->1
Simultaneous Mode 4 run on GOES-16 and GOES-17 12/04/2018 12/04/2018 Ran from 18 to 22 UTC; PDA output had some data gaps.
Mode 6J 24-hour test run 12/27/2018 12/28/2018 Run from 8 to 8 UTC. Navigation of visible and IR channels is in family with the previous versions of Mode 6. Time schedule appeared to be shorter than expected.
Mode 6L 24-hour test run 12/28/2018 12/29/2018 Run from 8 to 8 UTC. Navigation of visible channels is in family with the previous versions of Mode 6. Navigation of IR channels was broken.
Mode 6K 24-hour test run 12/29/2018 12/30/2018 Run from 8 to 8 UTC. Navigation of visible channels is in family with the previous versions of Mode 6. Navigation of IR channels was broken.
STAR SCDR outage 1/5/2019 1/7/2019 STAR SCDR did not accept G16/G17 data from PDA. The gap was later backfilled. Some backfill data were taken from GRB distribution. It was earlier noted by AWG that some GRB data may cause problems in L2 processing.
Mode 6M 48-hour test run 1/28/2019 1/30/2019 Run from 13 to 13 UTC; CWG evaluation of radiometric and geometric calibration is positive.
DO.07 has a code bug in implemetation of CDRL 80 Rev F 10/15/2018 04/08/2019 In equation 6.5-1 integration factor was erroneously divided twice during the Q-scaling in DO.07.( WR5407/WR6348).
GOES-17 operational declaration as GOES-West 02/12/2019 GOES-17 is officially declared as the GOES-West satellite at 137.2 degrees west.
Initial gainset switch time 7/26/2018 10/24/2018 Gainset switch time set at the beginning of the PLT/PLPT period. Gainset 1->3: 0600 UTC. Gainset 3->1: 1500 UTC.
Gainset swith time after the satellite drift 11/13/2018 12/06/2018 Gainset 1->3: 0900 UTC. Gainset 3->1: 1900 UTC.
Gainset 3 turned on 1/18/2019 Gainset 1->3: 0830UTC and gainset 3->1: 1930UTC.
Gainset 3 turned off 12/06/2018 1/18/2019 Gainset 3 was turned off during this period.
Gainset switch time 11/30/2018 12/06/2018 Gainset 1->3: from 0900 UTC to 0830 UTC. Gainset 3->1: from 1900 UTC to 1930 UTC
Mode 6 72-hour test run 2/19/2019 2/22/2019 Version 6M was run from 15 to 15 UTC.
Mode 6 56-hour test run 2/25/2019 2/27/2019 Version 6M was run from 14 to 22 UTC.
G17 B04 and B06 GS gain decreasing 8/02/2018 8/27/2018 G17 B04 and B06 solar calibration gain values became abnormally decreasing.Resulting in dark trending of L1B radiance during this period(ADR754).
G17 B04 and B06 GS gain decreasing 12/26/2018 2/25/2019 G17 B04 and B06 solar cal gains displayed abnormal decreasing trending again. It resulted in dark trending of L1B radiance at these two channels. The previous event was in August 2018. ADR 754 and ADR 909.
G17 Mode 6 test run 03/19/2019 03/22/2019 Version 6M was run from 14:00 to 14:00 UTC.
G17 IR bands gain set3 turn off 03/16/2019 G17 IR bands gain set3 turn off.
G17 BDS updates and DRL for selected IR detectors 03/22/2019 03/22/2019 G17 BDS map was updated for the following detectors to reduce striping and noise: B8D91 / B9D92 / B10D50 / B10D175 / B11D35 / B11D49 / B11D216 / B11D300 / B12D45 / B12D133 / B12D198 / B12D202 / B12D290 / B12D293 / B13D196 / B15D333 and 137. DRL was implemented for B16 Det#28
G17 B10 detector 231 saturation 03/21/2019 04/01/2019 The detector saturated and caused serious striping in L1b images. The striping was mitigated by manual memory flush executed at ~14z on 2019-03-22. The BDS map of the detector is updated on 04/01/2019 to resolve the anomaly.
G17 IR bands gain set III turned back on 03/28/2019 G17 IR bands gain set III was turned back on. Now the instrument is operated at the interleaved gain set I/III setting. Gainset I to III is switched at 0900UTC and Gainset III to I at 1900UTC.
G17 Yaw Flip 03/27/2019 03/27/2019 GOES-17 Yaw Flip accomplished on March 27 2019 2035-2159 UT; ABI data outages - incomplete last pre-flip images; few empty frames. Image navigation is good for all post-flip images.
G17 Transition to Mode 6 04/02/2019 New baseline Mode 6M is started at 16:00 UTC ; it replaced Mode 3C; link: https://www.ospo.noaa.gov/Operations/GOES/west/GOES-17_Scan_Mode_6.html.
G17 VNIR gain change with RevG algorithm 04/08/2019 G17 VNIR bands gains changed with the implementation of CDRL-80 RevG algorithm (DO.07.02) using the 04/02 solar cal data.
Abnormal G17 VNIR gains since the solar calibration event on April 12 2019 04/12/2019 G17 VNIR B04 and 06 experienced unexpected large drops of the operational solar cal gain values after the solar calibration event on 04/12/2019. Compared to the solar cal. gains on 04/08/2019 (DO.07.02) which were calculated based on the solar event on 04/02/2019. B04 dropped by 9.2%/B05 by 1.4%/B06 by 5.6%. All the changes were more than expected.
CPAF update 04/18/2019 04/18/2019 Band 7 resampler update at 18:32 UTC to correct the cold pixels around fire (CPAF).
G17 B02 bias correction 04/27/2019 G17 B02 bias correction was applied in GS at 15:40UTC on 04/27/2019 to reduce the radiance by about 6.9%. RevG algorithm was also implemented to cause the gain changes at B04-B06.
G17 Navigation anomaly 04/23/2019 4/23/2019 Anomalous navigation shifts were observed by CWG in the full disk image at 15:10 UTC for all channels. The images are shifted in both North-South and East-West directions. The largest shifts are in the EW direction of VNIR channels and amount up to 15 urad. The CONUS and MESO images produced in the period 15:10 to 15:20 may also be affected.
Kalman filter reset 04/27/2019 04/27/2019 Kalman filter reset after 13 UTC caused slightly elevated navigation residuals and ABI data gap from 1:30 to 3:20 UTC.
G17 Solar cal lunar intrusion rejection LUT update 05/05/2019 05/05/2019 Reduced striping or corrected bias at the G17 VNIR bands due to the update of solar spacelook count lunar intrusion rejection LUTs.
G17 Periodic striping for B11 04/23/2019 05/03/2019 Serious striping between 1900 - 0900 UTC daily caused by a detector (D241) saturation during Gain Set I. The detector was replaced at 1500 UTC on 05-03-2019.
Navigation anomaly 05/04/2019 05/04/2019 Anomalous navigation shifts and image loss were observed by CWG in the full disk image from 13:00 to 14:00 UTC for all channels. The images are shifted in both North-South and East-West directions. The largest shifts are in the EW direction of VNIR channels and amount up to 30 urad. The CONUS and MESO images produced in that period may also be affected.
Navigation anomaly 05/05/2019 05/05/2019 CWG has observed relatively-elevated EW navigation shift of up to 5 urad in WNIR channel full-disk images at 15:10 UTC. CONUS and MESO images collected from 15:10 to 15:20 UTC may also be affected.
Update the Kalman filter control LUTs 05/10/2019 05/10/2019 Deployed at 15:29 UTC this update is a temporary fix for the anomaly reported by ADR 955. This temporary fix will resolve the periodic Kalman filter issues that have been seen in the recent weeks and reported in ADR 955. The fix will be removed when the full fix is delivered in DO 09.
Gainset III removal 05/28/2019 07/15/2019 The daily gainset I and gainset III swap stopped on 05/28/2019.
North-South striping in GOES-17 Channel 2 image 05/31/2019 Multiple stripes along the north-south direction are found around 1830 - 1840 UTC since May 31 2019 (when clear). The stripes are centered at a solar farm in southern Calibration (35:20N/120:04W) and extend throughout one swath. They may persist until mid July.
Large image navigation errors 07/10/2019 07/10/2019 ABI images taken around 15:20 and 15:30 UTC may have large navigation errors due to the Kalman filter reset caused by the scheduled HBT flush.
Gainset III resume 07/16/2019 G17 IR bands gainset III daily turned on at 09:00Z; switched to gainset I at 19:00Z.
Sporadic missing data with DQF=-1 in CONUS/FD images 06/22/2019 06/22/2019 Missing data (radiance = 4095)with DQF=-1 was found in the CONUS/FD images at various bands and time that day. Similar incidents probably happened before.
Major GS upgrade to DO.08.00.02 07/25/2019 including the corrections of several issues concerning L1b products and the implementation of Predictive Calibration to improve GOES-17 Channel 8-16 calibration when the FPM temperature becomes unstable.
G17 ABI FD data gap 08/14/2019 08/14/2019 Data gap from 2019-08-14 0640 UTC to 2019-08-14 0830 UTC is caused by East-West encoder error or anomaly. It was resumed back to normal with the standard procedures.
G17 ABI FD data gap 08/15/2019 08/15/2019 Data gap from 2019-08-15 0540 UTC to 2019-08-15 1630 UTC is caused by encoder error or anomaly. It was resumed back to normal after parking the ABI sensor.
G17 ABI scanner calibration 08/16/2019 08/16/2019 ABI data outage from 2019-08-16 22:30 UTC to 2019-08-16 23:00 UTC due to the scanner calibration.
G17 ABI scanner calibration 08/17/2019 08/17/2019 ABI data outage from 2019-08-17 00:00 UTC to 2019-08-17 00:30 UTC due to the scanner verification.
G17 ABI navigation anomaly 07/31/2019 07/31/2019 Navigation shift is detected by CENRAIS in the full-disk image of 22:50 UTC in the EW and NS directions for all channels. In the channel 2 image the shift is about 8 microradians in the East direction and 15 microradians in the North direction. Similar shifts may be present in the other imagery taken between 22:50 and 23:00 UTC.
G17 band 11 striping 08/17/2019 09/09/2019 Channel 11 detector #49 (0-based index) spacelook count was unstable since 19:00z on 2019-08-17 and detector #305 was saturated since 02:15z on 2019-08-22. Both caused stripes in the Ch11 L1b image. The BDS of detector #49 was updated on 08/27 at 10:23z and the BDS of detector #305 was updated on 09/09 at 15:21z to remove the stripes.
Streaks at B02/B07/B16 images 06/27/2019 06/27/2019 Streaks can be observed at the G17 B02/B07/B16 FD/CONUS/MESO images collected at 05:12Z on 06/27/2019 due to the lunar intrusion in the spacelook.
G17 IR bands gain set3 turn off 09/10/2019 G17 IR bands gain set 3 is turned off.
G17 Yaw Flip 09/09/2019 09/09/2019 Yaw flip was carried out between 20:40 and 21:39 UTC. No data were provided during that period. Image navigation and data integrity after the flip are nominal.
Streaks at B02/B07/B10/B15 09/15/2019 09/15/2019 Streaks at the G17 B02/B07/B10/B15 FD/CONUS/MES images collected at 23:06 - 23:10Z.
G17 ABI image gap 09/20/2019 09/20/2019 17:20 and 17:30 UTC FD images are missing and 17:40 UTC FD images are partial for all channels. The data gap is caused by trial of antenna swap at NSOF. Only PDA output of FD and other images is affected
pCal on/off update 10/06/2019 The pCal is turned on at ~8:40z and off at ~19:20z so that the pCal on/off time will not collide with the gain set switch time at 08:30z and 19:30z.
G17 ABI TDQF LUT update 10/02/2019 10/02/2019 The G17 ABI Temperature Data Quality Flags (TDQF) thresholds for the IR bands were updated to flag the saturated data more effectively.
G17 ABI degraded images 10/01/2019 10/01/2019 ABI images dated from 13:39 to 22:30 UTC may be degraded due to the work on star tracker. Anomaly has impacted swath collection and alignment. Images may have missing blocks. Navigation errors are elevated at the start and the end of the period.
G17 IR bands gain set III resumed 10/06/2019 G17 IR bands gain set III was turned back on 10/06/2019. Now the instrument is operated at the interleaved gain set I/III setting. Gainset I to III is switched at 0830UTC and Gainset III to I at 1930UTC.
G17 cooling test I 10/15/2019 10/15/2019 The first phase test for the ABI 30 minute cooling timeline was conducted between 11:00 - 12:00 UTC. Only two Mode 4 timeline images were disseminated during this one hour. Some Band 2 data were missing.
Gain Optimization Test I 10/15/2019 10/15/2019 The gain adjustment test was conducted for B13/B14 during the Phase I cooling test period at approximately 11UTC-12UTC. The L1b images were severely damaged because the cooling test was executed concurrently. No noticeable impact in L1b was observed after 12 UTC.
G17 Cooling Test - Phase II 10/18/2019 10/21/2019 The second phase test for the ABI 30 minute cooling timeline was conducted between 0600 - 1200 UTC on 10/18/2019 - 10/21/2019. Only one Mode 4 timeline images were disseminated at every 30 minutes during these cooling test periods.
Gain Optimization Test II 10/18/2019 10/21/2019 The detector gains are adjusted for Bands 13 & 14 during the gain-set III period of 08:30 UTC - 19:30 UTC on 10/18/2019 - 10/21/2019. The adjustment adds the data availability for band 13 and reduces the noise for band 14.
Gain Optimization Test III 11/05/2019 11/06/2019 The detector gains are adjusted for Bands 13 & 14 during the gain-set III period of 08:30 UTC - 19:30 UTC on 11/05/2019 - 11/06/2019. Serious striping may be observed in Band 13 L1b images from 19:30 UTC to 19:35 UTC on 11/05/2019 and 11/06/2019 caused by the latched spacelook counts which were triggered by this gain optimization test.
Gain Optimization Test IV 11/07/2019 11/08/2019 The detector gains are adjusted for Band 14 during the gain-set period of 08:30 UTC - 19:30 UTC.
G17 B02 Q LUT Update 11/18/2019 11/18/2019 GOES-17 ABI B02 calibration quadratic term (Q) was updated with the CDRL-079 Rev T. This update only changed the Q values for some BDS detectors from Column 2. Users may not detect the change in the L1B data.
ABI scanner calibration 1 11/13/2019 11/13/2019 Due to the ABI scanner calibration the ABI L1b data were missing from 22:40 to 23:10 UTC.
ABI scanner calibration 2 11/14/2019 11/14/2019 Due to the ABI scanner calibration the ABI L1b data were missing from 00:00 to 00:20 UTC.
Streaks caused by lunar intrusion in space-looks 12/20/2019 12/20/2019 Streaks exist in the 04:40Z timeline L1b images for bands 10 and 10 that lasts 5 minutes and for bands 5-6 and 11-12 within one FD and one MESO scans. Both are caused by lunar intrusion into space-looks.
G17 IR bands gain set3 turn off 12/15/2019 12/15/2019 G17 IR bands gain set 3 is turned off.
G17 IR bands gain set3 turn on 12/19/2019 12/19/2019 G17 IR bands gain set 3 is turned on at 08:30 UTC.
Bands 12-16 anomalous images 12/31/2019 12/31/2019 The images of bands 12-16 of the first half of 19:30Z timeline are anomalous because the detector gain calculation is incorrect.
B09 BDS update 01/09/2020 01/09/2020 B09 Det#36 (indexing from 0) BDS was updated at 17:31 UTC on 01/09/2020. This is aimed to resolve the intermittent striping caused by the unstable detector responsivity starting on 01/06/2020.
B10 BDS update 02/07/2020 02/07/2020 B10 Det#297 (indexing from 0) BDS was updated at ~18:23Z on 02/07/2020 to resolve the moderate striping caused by the unstable detector responsibility starting on 02/03/2020.
G17 ABI Cooling Timeline Engineering Test 02/19/2020 02/19/2020 GOES-17 ABI experienced two hours (1800-2000 UTC) of a cooling timeline engineering test using the "Flex Mode 3" on these two days.
Mode 3 cooling timeline implementation 02/26/2020 03/01/2020 The Mode 3 cooling timeline ("Flex Mode 3") was implemented for 6 hours from 0600 UTC to 1200 UTC from Feb 26
pCal anomaly 03/03/2020 03/05/2020 (1) The predictive calibration (pCal) algorithm for GOES-17 ABI bands 8-16 was not turned on during the unstable IR FPM temperature time on March 3-4 2020 resulting in relatively large radiometric calibration uncertainty for the data of these bands collected during this period; (2) the pCal algorithm was incorrectly implemented and collided with the gain-set switch events at 19:30 UTC on March 4 and 08:30 UTC on March 5 resulting in degraded images of bands 8-16 at 19:30-19:40UTC on March 4 and 08:30-08:40 UTC on March 5.
B11 BDS update 03/10/2020 03/10/2020 G17 Band11 Detector#081 BDS was updated at ~16:58 UTC on March 10 2020 to resolve an intermitent striping starting since late February 2020.
Real-time op. instr. cal files outage 02/23/2020 The real-time operational instrument calibration files are not available via PDA.
Gainset switch turn-off 03/18/2020 03/24/2020 The gainset switch was disabled at 08:30 UTC on 03/18/2020. It turned back on 03/25/2020.
G17 HBT flush 03/10/2020 03/10/2020 GOES-17 HBT flush was performed from 15:10 to 15:40 UTC
G17 Yaw-flip 04/06/2020 04/06/2020 G17 yaw flip was carried out on 04/06/2020. Data outage via PDA was observed between 20:40-21:33UTC. Blank or partial images can be observed during that period. Image navigation is nominal and the pixel quality flag is correct after the flip. More details in CWG presentations by T. Schmit and S. Guo on 2020-04-08.
Change in the Gainset Switch Time 04/07/2020 05/31/2020 The Gain Set III has been in use between 0900-1900 UTC since 7 April 2020
G17 B12 Det159 Anomaly 04/13/2020 04/16/2020 G17 B12 Det#159 became saturated at ~16:25Z on 04/13/2020; causing strong striping at the B12 images. The striping was reduced at ~17:54Z on 04/14/2020 by a change in the resampling processing. The BDS update occurred at ~16:56Z on 04/16/2020 to resolve the anomaly. Details in the OSPO notification Issued on Apr 14/2020 1840Z and the CWG meeting reports on April 15 and April 22/ 2020.
pCal update 04/28/2020 The pCal control was updated with ADR1001 Case10 on April 28/2020. It is scheduled to function till August 2020
Mode 3 Cooling Timeline 04/09/2020 05/01/2020 The Mode 3 cooling timeline ("Flex Mode 3") was implemented to extend the valid ABI data by reducing the peak IR FPM temperature. It was executed from 0600-1200 UTC on 04/09/2020 - 04/24/2020; from 0600-1130 UTC on 04/25/2020; and from 0600 - 1200 UTC on 04/26/2020 to 05/01/2020. Details in the OSPO notification issued on 04/06/2020 at 1310Z and the reports in the CWG meeting on 04/29/2020 and 05/06/2020.
B09 Det193 Anomaly 05/17/2020 05/21/2020 G17 B09 Det#193 became unstable at ~14:00Z on 05/17/2020 and resulted in striping images since then. The striping issue was resolved with the BDS update at ~15:24Z on 05/21/2020. Details on the reports in the CWG meeting on 05/27/2020.
GOES-17 INR anomaly 05/04/2020 05/04/2020 Large East-West and North-South image navigation errors in the 13:00 to 13:30 UTC FD images for reflective channels (other formats may also be affected) due to reset of the Kalman filter. Channel errors are up to 12 urad. Details are in the presentation by S. Guo and V. Kondratovich at the 2020-05-27 CWG meeting.
Gainset I Only 06/01/2020 06/02/2020 The gainset III was turned off
Gainset III turn-on 06/03/2020 06/06/2020 The gainset III calibration configuration was turned on during these four days.
Gainset III turn-off 06/07/2020 Only the gainset I calibration configuration was applied
G17 B09 Det028 Anomaly 05/31/2020 06/04/2020 G17 B09 Det028 became unstable at ~10:09Z on 05/31/2020 resulting in intermittent striping. The striping issue was resolved with the BDS update on 06/04/2020. Details reported in the CWG meetings on 06/03/2020 and 06/10/2020.
G17 B11 Det300 Anomaly 06/03/2020 06/04/2020 Saturation response occurred at ~19:00Z on 06/03/2020. The strong striping was reduced with the manual trigger of the "band-aid" algorithm. A BDS update was implemented on 06/04/2020 to replace the anomalous detector. Details reported in the CWG meeting on 06/10/2020.
Gainset III resume 07/03/2020 Gainset switch returned on 07/03/2020.
B08 BDS update 07/06/2020 07/06/2020 The BDS of G17 B08 Det#84 was updated on 07/06/2020 at ~20:20:00 UTC. This is to resolve the striping caused by unstable detector responsivity since 06/29/2020. Details are in the presentation by Paul Douglas at the 2020-07-08 CWG meeting.
GOES-17 INR anomaly 06/22/2020 06/22/2020 Large East-West IR image navigation errors (up to 30 urad) at 19:30 UTC were detected by CWG in PDA output. Details are in the presentation by S. Guo and V. Kondratovich at the 2020-06-30 CWG meeting.
GOES-17 ABI Data Loss 07/01/2020 07/01/2020 ABI data from 16:50 to 17:10 UTC were lost because of the signal reception problem. Image navigation is strongly affected between 17:10 and 17:20
GOES-17 ABI INR Anomaly 07/04/2020 07/04/2020 Large navigation errors (up to 8 urad) in reflective channels at 14:30 UTC. Details are in the presentation by S. Guo and V. Kondratovich at the 2020-07-08 CWG meeting.
Banding and Striping Images 07/12/2020 07/12/2020 The Moon appeared in the G17 ICT spacelook positions at ~03:15z and ~03:20z. These intrusions caused the strong bandings at the B09 and B13 images between 03:15-03:20z and at the B10 and B15 images between 03:20z - 03:25z. One strong banding can also be observed at the B04 FD swath and the MESO image at ~03:15z. The strong striping at the B13 images since 03:20z was caused by the latched spacelook and was resolved by the manual reset of the spacelook at ~04:14z. Details on these anomalies are available in the report by Yu et al. in the CWG weekly meeting on 07/15/2020.
B11 stripings 08/06/2020 08/21/2020 The G17 B11 images experienced an intermittent striping between ~12:51 UTC on 08/06/2020 and ~21:02 UTC on 08/18/2020 and a strong striping between ~21:02 UTC and ~23:55 UTC on 08/18/2020. The intermittent striping was caused by the unstable detector at the detector row #213 and it was resolved by a BDS update at ~21:02 UTC on 08/18/2020. The strong striping caused an inadvertent BDS update at Row#300 at ~21:02z; and mitigated with a manual spacelook reset at ~23:55z on 08/18/2020 to trigger the band-aid remedy; and recovered with the reverted BDS update at ~18:28 UTC on 08/21/2020. Details in the reports at the CWG 08/19/2020 meeting.
B08 striping and BDS update 08/15/2020 08/24/2020 G17 B08 experienced an intermittent striping starting at ~22:42 UTC on 08/15/2020. The issue was resolved with the BDS update for the detector at Row#198 at ~17:30 UTC on 08/24/2020.
Cooling timeline implementation 08/11/2020 09/01/2020 he Mode3 cooling timeline was implemented between 06:00 - 12:00 UTC everyday from 08/11/2020 to 09/01/2020. This timeline is aimed to gain more valid earth scenes around the satellite midnight by reducing the FPM temperature in the warm time of the day in the eclipse season. When this cooling timeline is executed only one FD and 15 MESO images are collected in each timeline. Details can be found in the CWG meeting presentations on 09/02/2020.
Gainset switch time update 09/09/2020 The gainset switch times are changed from 09:00z/19:00z to 08:30z/19:30z
GOES-17 Yaw Flip 09/08/2020 09/08/2020 G17 Yaw Flip was performed between 20:40-21:39 UTC (no image in this period). All images before and after the yaw flip are nominal. No impact on INR. Details are in the presentation by S. Guo and V. Kondratovich at the 2020-09-09 CWG meeting.
Gainset I only 09/13/2020 10/02/2020 Gainset I only during this period. Validation reported in CWG meetings on 09/16/2020 and 10/07/2020.
B10 striping & BDS update 09/25/2020 09/30/2020 G17 B10 images experienced weak to moderate striping since 09/25/2020 due the unstable responsivity and elevated noise at Det#175. The issue was resolved with the BDS update on 09/30/2020. Detailed information is available at the report in the CWG meeting on 09/30/2020.
Gainset III not switched on 10/11/2020 10/11/2020 The Gainset III was not turned on this day.It caused unexpected striping or even invalid images at some IR bands near the satellite midnight time. Details available at the CWG meeting on 10/14/2020.
Cooling Operation 10/14/2020 10/31/2020 The operational cooling timeline was implemented between 06:00-12:00 UTC everyday from 10/14/2020 to 10/31/2020. This cooling operation is aimed to gain more valid earth scenes around the satellite midnight time. More details are available in the report in the CWG meeting on 11/04/2020.
Corrupted IR images at 8:30 UTC 11/06/2020 11/10/2020 Corrupted (degraded) B08-B16 images can be observed at 08:30 - 08:40Z from 11/06/2020-11/10/2020. This was due to the collision of the pCal algorithm with the gainset switch. The anomaly was corrected on 11/11/2020 with the update of pCal LUT. Details can be found at the ABI status report in the CWG meeting on 11/18/2020.
ABI Encoder Calibration 11/17/2020 11/17/2020 ABI Encoder Calibration was performed between 22:30-23:00 UTC. No images were produced during the period. No impact to the INR after the period.
ABI Encoder Calibration 11/18/2020 11/18/2020 ABI Encoder Calibration was performed between 00:00-00:30 UTC. No images were produced during the period. No impact to the INR after the period.
Calibration Variation at ~09:00Z 11/11/2020 11/25/2020 A small calibration variation (<~0.4K) can be observed at several G17 IR bands around 09:10 - 09:30Z from 11/11/2020 - 11/25/2020. It was caused by a later turn-on of the pCal algorithm with pCal LUT Case11. The issue was resolved with the deployment of a typo corrected LUT on November 25. More details in the CWG meeting on 12/02/2020.
stripes at B08 B11 and B12 01/06/2021 01/06/2021 Stripes can be observed at the B11 and B12 images from ~03:40 - ~03:50 UTC and the B08 images from ~03:45-03:50 UTC on 01/06/2021. The anomaly was caused by the lunar intrusion into the ICT spacelook observations for some detectors of these three bands. Reference can be found at the CWG meeting on 01/06/2021.
B09 striping and BDS update 01/24/2021 01/29/2021 G17 B09 experienced strong striping related to a saturated detector from ~22:30 UTC to ~23:55 UTC on 01/24/2021 and 15:40 - 15:45 UTC on 01/29/2021. The striping was mitigated by manually forcing the detector observations to skip the resampling algorithm at ~23:55 UTC on 01/24/2021 and was then resolved with the BDS update at ~15:40 UTC on 01/29/2021. Strong striping could be observed from 15:40-15:45UTC after the BDS update and then disappeared after 15:45 UTC. More details available in the CWG meeting slides presented on 02/03/2021.
Cooling operation 02/06/2021 02/28/2021 The flexible M3 cooling timeline was applied from 06:00-12:00Z every day between 02/06/2021 and 02/28/2021. The cooling timeline was used to reduce the IR FPM temperature at night time and thus gained some valid images in the otherwise saturation period.
Calibration Anomaly at 09:00-10:00Z 02/27/2021 03/03/2021 Abnormal calibration variation was observed at some G17 IR channels at ~9-10Z from 02/27/2021 to 03/03/2021. The anomaly was associated with the small bump of G17 focal plane module (FPM) temperature during this time period with the maximum magnitude occurred on Mar. 1
G17 calibration anomaly at Bands 08-16 03/08/2021 03/08/2021 The pCal algorithm was not deployed on 03/08/2021 and this led to large radiometric calibration error for Bands 08-16 during the satellite night time when the focal plane module temperature was floating. The pCal algorithm was resumed on 03/09/2021. Details in the CWG meeting reports on 03/10/2021.
Corrupted images caused by GSS and pCal collision 03/31/2021 03/31/2021 The images of Ch08-16 collected at 19:30-19:35 UTC were corrupted due to the pCal and gainset switch collision.
Stripes caused by Lunar Intrusions 03/28/2021 03/28/2021 Stripes can be observed in the images of some G17 channels collected at 20:30-20:50UTC and 22:10-22:20 UTC. These stripes are caused by the lunar intrusions into the spacelook positions. More information was available in the CWG meeting on 03/31/2021 and 04/07/2021.
VNIR channel gains update via BYOG 04/05/2021 04/05/2021 GOES-17 VNIR channel gains were implemented into operation via Bright-Your-Own-Gain (BYOG) at 17:31 UTC on 5 April 2021. It was about 10 days after the scheduled solar calibration event conducted on 03/26/2021. The effort was to minimize the impact of abnormal PG solar calibration processing. Detailed information can be found in the CWG meeting on 03/31/2021 and 04/07/2021.
B12 Striping & BDS update 04/02/2021 04/08/2021 G17 B12 experienced intermittent striping due to the underperformance of Det#56 since 04/02/2021. This issue was resolved with the BDS update on 04/08/2021. Details available at CWG meeting on 04/07/2021 and 04/14/2021.
B08 Striping and BDS update 12/12/2020 05/14/2021 Intermittent striping appeared in the G17 B08 images around mid-December 2020 and disappeared on 05/14/2021 after the BDS update for Det#165. Details in the CWG presentations on 05/12/2021 and 05/19/2021.
B12 Striping and BDS update 05/04/2021 05/14/2021 Striping appeared in the G17 B12 images on 05/04/2021. It disappeared on 05/14/2021 after the BDS update for Det#207. Details in the CWG presentations on 05/12/2021 and 05/19/2021.
B13 BDS update 6/11/2021 6/11/2021 The BDS of B13 Detector row #280 was updated at ~16:05 UTC on 06/11/2021 to improve the image quality associated with the elevated noise of this detector since 05/29/2021. Details in the CWG presentations on 06/02/2021 and 06/16/2021.
ABI image sequential anomaly 05/29/2021 05/29/2021 ABI L1b all channel images were not distributed between 20:59-21:35 UTC. Blank FD images were distributed at 21:30. Refer to ESPC Notification and S. Guo and V. Kondratovich CWG meeting presentation on 6/02/2021.
B12 striping and BDS update 06/24/2021 07/02/2021 The images of B12 displayed striping caused by the elevated noise of Det#313 since 06/24/2021. The issue was improved with the BDS map update on 07/02/2021. Details are available in the CWG presentations on 06/30/2021 and 07/07/2021.
Broad striping by lunar intrusion 07/01/2021 07/01/2021 Broad striping can be observed in the B05/B10/B11/B12 images at ~03:27 UTC on 07/01/2021. It was caused by a lunar intrusion into the spacelook scene. Detailed information is available in the CWG presentation on 07/07/2021.
Satellite Anomaly and ABI reset 07/09/2021 07/09/2021 GOES-17 satellite experienced an anomaly on 07/09/2021 in resulting in a data outage from 06:48 - 12:20 UTC. The ABI instrument was reset during this period. Details are available in the CWG presentations on 07/14/2021.
B12 striping and the BDS map restoration 07/09/2021 07/12/2021 B12 images showed striping after the instrument reset on 07/09/2021. The issue was resolved with the restoration of the B12 BDS map for Det#313 on 07/12/2021. Details are available in the CWG presentations on 07/14/2021.
B09 Transient Striping 07/12/2021 07/12/2021 B09 images showed a transient striping at ~16:39Z. Details are available in the CWG presentations on 07/14/2021.
Satellite Anomaly and ABI Reset 07/22/2021 07/23/2021 G17 satellite entered into the safehold mode at ~05:37Z on 07/22/2021. ABI was reset and the L1b data was recovered via PDA at ~12:40Z on 07/23/2021. Impacts of this reset on the L1b data are available in the CWG meeting slides on 07/28/2021.
Broad striping by lunar intrusion 07/30/2021 07/30/2021 Broad striping can be observed in the B05/B10/B15 images at ~02:40 UTC on 07/30/2021. It was caused by a lunar intrusion into the spacelook scene. Detailed information is available in the CWG presentation on 08/04/2021.
Broad striping in FD images by lunar intrusion 07/31/2021 07/31/2021 Broad striping can be observed in FD only in the B04/B09/B13 images at ~03:00 UTC on 07/31/2021. It was caused by a lunar intrusion into the spacelook scene. Their MESO/CONUS images were normal. The B03 spacelook was slightly contaminated and the striping signal is weak and not visible. Detailed information is available in the CWG presentation on 08/04/2021.
Broad striping by lunar intrusion 07/28/2021 07/28/2021 Broad striping can be observed in the B03/B08/B09/B13/B14 FD images at 01:00 UTC (CONUS at 01:006 UTC)(MESO2 at 01:08UTC) on 07/28/2021. It was caused by a lunar intrusion into the spacelook scene. Detailed information is available in the CWG presentation on 08/04/2021.
G17 IR Calibration Anomaly 08/06/2021 08/06/2021 G17 ABI experienced a series of anomalies in the images collected at ~09:00 - 09:15Z and ~12Z on 08/06/2021. The root cause of these one-day event anomalies was a large slip of the beginning time of the M3 timelines that day. At ~9UTC it caused the gainset switch occurring within the timeline. At ~12UTC incomplete timeline data were collected. Details are available in the slides presented in the CWG meeting on 08/11/2021 and 08/18/2021.
G17 IR calibration anomaly 06/18/2021 07/29/2021 G17 ABI B08-B16 experienced a radiometric calibration anomaly at several timelines when the FPM temperature started to increase at night since 06/18/2021. This led to a bump feature in the GEO-GEO inter-comparison. The issue was resolved with the update of the pCal LUT on 07/29/2021. Details are available in the CWG meeting report on 08/04/2021.
B09 striping and BDS update 08/03/2021 08/19/2021 Striping was occasionally observed in B09 at 06:30z since 08/03. It was caused by the elevated noise of the Det#98. The issue was improved with the BDS map update on 08/19/2021. Details are available in the CWG presentations on 08/11/2021 and 08/25/2021.
Large solar calibration variation 08/07/2021 The G17 B03 radiance may have been lower by up to 3.7% (1.6% for B02) since 08/07/2021 when a solar calibration on that day occurred at high FPM temperature. This may end after the next scheduled solar calibration on Sept 8. More details were reported to the CWG meeting on 08/11/2021.
GOES-17 Yaw Flip 09/07/2021 09/07/2021 G17 Yaw Flip was performed between 20:40-21:29 UTC (no image in this period). All images (FD
Broad striping in FD/CONUS images by lunar intrusion 08/24/2021 08/24/2021 road striping was observed in the B10/B11/B12/B15 FD L1b images at ~23:50 UTC and in their CONUS L1b images at ~23:56z on 08/24/2021. It was caused by the lunar intrusion into the spacelook scene. Detailed information is available in the CWG presentation on 09/08/2021.
Broad striping in FD images by lunar intrusion 08/27/2021 08/27/2021 Broad striping was observed in the B09/B13 FD L1b images at ~01:00 UTC and in B10/B11(very faint)/B12(faint)/B15 FD L1b images at ~01:10 UTC on 08/27/2021. It was caused by the lunar intrusion into the spacelook scene. Detailed information is available in the CWG presentation on 09/08/2021.
BOA Anomaly 09/03/2021 09/03/2021 G17 Ch08-16 FD images were corrupted in the bright object avoidance (BOA) swaths at 09:00Z on 09/03/2021 and 09/04/2021. The root cause was the early midnight spacelook swap on these two days. Details are available in the CWG meeting slides on 09/15/2021.
Cooling operation 08/06/2021 09/07/2021 The flexible M3 cooling timeline was applied from 06:00-12:00Z every day between 08/06/2021 and 09/07/2021. The cooling timeline was accidentally not implemented on 09/02/2021. Details in the CWG meeting reports on 09/15/2021 and 09/22/2021.
Broad striping in L1b images by a lunar intrusion 09/22/2021 09/22/2021 Broad striping was observed in the B08/B09/B13/B14 L1b images at 23:10 - 23:20 UTC on 09/22/2021. It was caused by the lunar intrusion into the SPL scene. Detailed information is available in the CWG slides presented by Qian on 09/29/2021.
Broad striping in L1b images by a lunar intrusion 09/21/2021 09/21/2021 G17 B08/B09/B10/B13/B14/B15 L1b images displayed a series of broad striping in the L1b images collected between 20:30 - 20:39 UTC on 09/21/2021. They were caused by the lunar intrusion in the nominal SPL and/or the ICT SPL. Detailed information is available in the CWG presentation by Qian on 09/29/2021.
Broad striping in L1b images by a lunar intrusion 09/20/2021 09/20/2021 Broad striping was observed in the B08/B09/B13/B14 FD L1b images at ~20:20 UTC and the CONUS L1b images at ~20:25 UTC and MESO1 at 20:28 UTC on 09/20/2021. It was caused by the lunar intrusion into the spacelook scene. Detailed information is available in the CWG presentation on 09/29/2021.
FPM setpoint temperature test - Part I 09/17/2021 09/20/2021 G17 ABI FPM setpoint temperature was tested at 78K during this period. Details in the CWG meetings between 09/27/2021 - 10/20/2021.
FPM setpoint temperature test - Part II 09/22/2021 09/24/2021 G17 ABI FPM setpoint temperature was tested at 84K during this period. Details in the CWG meetings between 09/27/2021 - 10/20/2021.
FPM setpoint temperature change 10/08/2021 G17 ABI FPM setpoint temperature has changed from 81K to 84K on 10/08/2021. Details in the CWG meetings on 10/06/2021 and 10/13/2021.
pCal LUT (Case 13) implemented 10/15/2021 The pCal control was updated with Case13 on 10/15/2021. Detailed information is available in the CWG presentation on 10/20/2021.
FPM setpoint test - Part III 10/18/2021 10/20/2021 G17 ABI FPM setpoint temperature was set at 87K during this period. Details in the CWG meeting on 10/20/2021.
Broad striping caused by lunar intrusion 10/17/2021 10/20/2021 Broad striping was observed due to lunar intrusion into the spacelook scene. The contaminated spacelook observations occurred at the following times for the specified bands and caused degradation in the products scanned during that time: 10/17 1829Z bands 08 11 12 14 16; 10/18 1856Z bands 11 12; 10/19 1943Z bands 08 10 15 16; 10/19 2113Z bands 09 13; 10/20 2040Z bands 08 14 16; 10/20 2130Z bands 08 16. Details in the CWG meeting on 10/27/2021.
FPM setpoint test - Part III (Extended) 10/25/2021 10/29/2021 G17 AB12/22/2021I FPM setpoint temperature was set at 87K during this period. Details in the CWG meeting on 11/03/2021.
Broad striping caused by lunar intrusion 11/13/2021 11/16/2021 Broad striping was observed due to lunar intrusion into the spacelook scene. The contaminated spacelook observations occurred at the following times for the specified bands and caused degradation in the products scanned during that time: 11/13 1640Z timeline in bands 8 and 14; 11/14 1650Z timeline in bands 11 and 12; 11/15 1740Z timeline in bands 9 and 13; 11/16 1830Z timeline in bands 8 11 12 and 16.
Broad striping by lunar intrusion 11/30/2021 11/30/2021 A broad striping was observed in the G17 ABI B10/B15 L1b images at 06:00 UTC on 11/30/2021 due to a lunar intrusion. Details are available in the CWG meeting report on 12/08/2021.
G17 IR calibration anomaly 11/30/2021 11/30/2021 The pCal algorithm was temporarily turned off on 11/30/2021 at 17:00 UTC after a short time data outage during 16:25 - 17:00 UTC. This was due to a ground system hardware issue. This affected radiometric calibration uncertainty for Bands 08-16 and the FPM temperature in the metadata at this timeline was abnormal.
BDS maps update 12/22/2021 12/22/2021 G17 BDS maps for B08-16 were updated at ~16:11Z on 12/22/2021 to improve the image quality of these IR channels. Details in the CWG meetings on 12/15/2021 and 12/22/2021.
Broad striping caused by lunar intrusion 12/12/2021 12/13/2021 Broad striping was observed due to lunar intrusion into the spacelook scene. The contaminated spacelook observations occurred at the following times for the specified bands and caused degradation in the products scanned during that time: 12/12 1530Z timeline in bands 11 and 12; 12/12 1540Z timeline in bands 08 11 12 14; 12/13 1620Z timeline in bands 11 and 12; 12/13 1630Z timeline in bands 09 and 13. Details in the CWG meeting on 12/22/2021.
Broad striping caused by lunar intrusion 12/26/2021 12/27/2021 Broad striping was observed due to lunar intrusion into the spacelook scene. The contaminated spacelook observations occurred at the following times for the specified bands and caused degradation in the products scanned during that time: 12/26 0300Z timeline in bands 10 and 15; 12/27 0400Z timeline in bands 09 and 13. Details in the CWG meeting on 01/05/2022.
Broad striping caused by lunar intrusion 01/06/2022 01/10/2022 Broad striping was observed due to lunar intrusion into the spacelook scene. The contaminated spacelook observations occurred at the following times for the specified bands and caused degradation in the products scanned during that time: 1/7 1300Z timeline in bands 08 and 14; 1/8 1330Z timeline in band 16; 1/9 at 1422Z in bands 08 09 13 and 14; and 1/10 in bands 09 10 13 15 at 1520Z.